# Recursive Neyman Algorithm for Optimum Sample Allocation under Box Constraints on Sample Sizes in Strata

### v3

@@ 2. CASE II:
    Take $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ with *$\mathcal{L}^{*},\,\mathcal{U}^{*}$ satisfying (3.4) and (3.5). Then, condition (B.5) becomes (3.5), while (B.6) is trivially met due to (3.2). Assume that $\mathcal{L}^{*}\neq\emptyset$ and $\mathcal{U}^{*}\neq\emptyset$ (for empty $\mathcal{L}^{*}$ or $\mathcal{U}^{*}$, (B.4), (B.7) and (B.8) are trivially met). Take an arbitrary $\tilde{s}>0$ such that
@@ C.1 Notation
$*.\mathcal{L}_{r}=\mathcal{L}_{r-1}\cup\widetilde{\mathcal{L}}_{r-1}=\bigcup\limits_{i=1}^{r-1}\widetilde{\mathcal{L}}_{i},\mathrlap{\qquad r=2,\ldots,r^{*}.}$ (C.1)
@@ C.2 Auxiliary remarks and lemmas
[…]
@@
if and only if $\mathbf{x}^{*}$ (computed at Step 6 of RNABOX algorithm) is: a regular allocation or a vertex allocation with $\mathcal{L}_{r^{*}}=\mathcal{H}$. Moreover, for $r^{*}\geq 1$,
~$*.\widetilde{\mathcal{L}}_{r}=\{h\in\mathcal{H}\setminus(\mathcal{L}_{r}\cup\mathcal{U}_{r}):\,A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})\leq m_{h}\},\mathrlap{\qquad r=1,\ldots,r^{*}.}$ (C.15)
@@
Let $r^{*}\geq 2$ and $r=1,\ldots,r^{*}-1$. Then, by (C.6), $\mathcal{L}_{r}\cup\mathcal{U}_{r}\subsetneq\mathcal{H}$. Following (C.13), the domain of discourse for $\mathcal{U}_{r}$ is $\mathcal{H}\setminus\mathcal{L}_{r}$, and in fact it is $\mathcal{H}\setminus(\mathcal{L}_{r}\cup\widetilde{\mathcal{L}}_{r})=\mathcal{H}\setminus\mathcal{L}_{r+1}$, since $\mathcal{U}_{r}\not\subset\widetilde{\mathcal{L}}_{r}$ as ensured by Step 4 of RNABOX. That is, both $\mathcal{U}_{r}$ and $\mathcal{U}_{r+1}$ have essentially the same domain of discourse, which is $\mathcal{H}\setminus\mathcal{L}_{r+1}$. Given this fact and the form of the set-builder predicate in (C.13) - (C.14) as well as equality $\mathcal{U}_{r^{*}}=\mathcal{H}\setminus\mathcal{L}_{r^{*}}$ for the case when $\mathbf{x}^{*}$ is a vertex allocation (for which (C.14) does not apply), we conclude that only one of the following two distinct cases is possible: $\mathcal{U}_{r}\supseteq\mathcal{U}_{r+1}$ or $\mathcal{U}_{r}\subsetneq\mathcal{U}_{r+1}$. The proof is by contradiction, that is, assume that (C.16) does not hold. Therefore, in view of the above observation, there exists $r\in\{1,\ldots,r^{*}-1\}$ such that $\mathcal{U}_{r}\subsetneq\mathcal{U}_{r+1}$. Then,
@@ 2. Vector $\mathbf{x}^{*}$ is a vertex allocation:
Hence (C.28) is proved. If $\mathbf{x}^{*}$ is a regular allocation, in view of Remark C.1, the same reasoning leading to inequality (C.35) clearly remains valid for $r=r^{*}-1,\,r^{*}\geq 2$. ∎
@@
    where the fact that above inequality is met for $h\in\mathcal{L}_{r^{*}-1}$ follows from (C.28). By comparing (C.45) and (C.46) we clearly see that (C.38) is satisfied. Lastly, equation (C.39) is fulfilled due to
@@
    where first equality follows from (C.1) while the second one follows from Step 3 of the RNABOX after referring to (D.3) and using the fact that $\mathcal{U}_{r^{*}}=\mathcal{H}\setminus\mathcal{L}_{r^{*}}$ for a vertex allocation.
@@
[…]
@@ Appendix E Appendix: Convex optimization scheme and the KKT conditions
where $\mathcal{D}\subseteq{\mathbb{R}}^{p},\,p\in{\mathbb{N}}_{+}$, is the optimization variable, the objective function $f:\mathcal{D}_{f}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}}$ and inequality constraint functions $g_{j}:\mathcal{D}_{g_{j}}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}},\,j=1,\ldots,\ell$, are convex, whilst equality constraint functions $w_{i}:\mathcal{D}_{w_{i}}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}},\,i=1,\ldots,k$, are affine. Here, $\mathcal{D}=\mathcal{D}_{f}\cap\bigcap_{i=1}^{k}\mathcal{D}_{w_{i}}\cap\bigcap_{j=1}^{\ell}\mathcal{D}_{g_{j}}$ denotes a common domain of all the functions. Point $\mathbf{x}\in\mathcal{D}$ is called feasible if it satisfies all of the constraints, otherwise the point is called infeasible. An optimization problem is called feasible if there exists $\mathbf{x}\in\mathcal{D}$ that is feasible, otherwise the problem is called infeasible.
@@
A point $\mathbf{x}^{*}\in\mathcal{D}\subseteq{\mathbb{R}}^{p}$ is a solution to the convex optimization problem (E.1) in which functions $g_{j},\,j=1,\ldots,\ell$, are affine if and only if there exist numbers $\lambda_{i}\in{\mathbb{R}}$, $i=1,\ldots,k$, and $\mu_{j}\geq 0$, $j=1,\ldots,\ell$, called KKT multipliers, such that
$\begin{gathered}\nabla f(\mathbf{x}^{*})+\sum_{i=1}^{k}\lambda_{i}\nabla w_{i}(\mathbf{x}^{*})+\sum_{j=1}^{\ell}\mu_{j}\nabla g_{j}(\mathbf{x}^{*})=\mathbf{0}\\ w_{i}(\mathbf{x}^{*})=0,\mathrlap{\qquad i=1,\ldots,k}\\ g_{j}(\mathbf{x}^{*})\leq 0\mathrlap{\qquad j=1,\ldots,\ell}\\ \mu_{j}g_{j}(\mathbf{x}^{*})=0\mathrlap{\qquad j=1,\ldots,\ell.}\end{gathered}$ (E.2)

### v4

@@ 2. CASE II:
    Take $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ with $\mathcal{L}^{*},\,\mathcal{U}^{*}$ satisfying (3.4) and (3.5). Then, condition (B.5) becomes (3.5), while (B.6) is trivially met due to (3.2). Assume that $\mathcal{L}^{*}\neq\emptyset$ and $\mathcal{U}^{*}\neq\emptyset$ (for empty $\mathcal{L}^{*}$ or $\mathcal{U}^{*}$, (B.4), (B.7) and (B.8) are trivially met). Take an arbitrary $\tilde{s}>0$ such that
@@ C.1 Notation
$\mathcal{L}_{r}=\mathcal{L}_{r-1}\cup\widetilde{\mathcal{L}}_{r-1}=\bigcup\limits_{i=1}^{r-1}\widetilde{\mathcal{L}}_{i},\mathrlap{\qquad r=2,\ldots,r^{*}.}$ (C.1)
@@ C.2 Auxiliary remarks and lemmas
$x^{**}_{h}=\begin{cases}M_{h},&\qquad{h\in\mathcal{U}_{r}\subseteq\mathcal{H}\setminus\mathcal{L}_{r}}\\ A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})<M_{h},&\mathrlap{\qquad h\in\mathcal{H}\setminus(\mathcal{L}_{r}\cup\mathcal{U}_{r}),}\end{cases}$ (C.12)
@@
if and only if $\mathbf{x}^{*}$ (computed at Step 6 of RNABOX algorithm) is: a regular allocation or a vertex allocation with $\mathcal{L}_{r^{*}}=\mathcal{H}$.
Moreover, for $r^{*}\geq 1$,
~$\widetilde{\mathcal{L}}_{r}=\{h\in\mathcal{H}\setminus(\mathcal{L}_{r}\cup\mathcal{U}_{r}):\,A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})\leq m_{h}\},\mathrlap{\qquad r=1,\ldots,r^{*}.}$ (C.15)
@@
Let $r^{*}\geq 2$ and $r=1,\ldots,r^{*}-1$. Then, by (C.6), $\mathcal{L}_{r}\cup\mathcal{U}_{r}\subsetneq\mathcal{H}$. Following (C.13), the domain of discourse for $\mathcal{U}_{r}$ is $\mathcal{H}\setminus\mathcal{L}_{r}$, and in fact it is $\mathcal{H}\setminus(\mathcal{L}_{r}\cup\widetilde{\mathcal{L}}_{r})=\mathcal{H}\setminus\mathcal{L}_{r+1}$, since $\mathcal{U}_{r}\not\subset\widetilde{\mathcal{L}}_{r}$ as ensured by Step 4 of RNABOX. That is, both $\mathcal{U}_{r}$ and $\mathcal{U}_{r+1}$ have essentially the same domain of discourse, which is $\mathcal{H}\setminus\mathcal{L}_{r+1}$. Given this fact and the form of the set-builder predicate in (C.13) - (C.14) as well as equality $\mathcal{U}_{r^{*}}=\mathcal{H}\setminus\mathcal{L}_{r^{*}}$ for the case when $\mathbf{x}^{*}$ is a vertex allocation (for which (C.14) does not apply), we conclude that only one of the following two distinct cases is possible: $\mathcal{U}_{r}\supseteq\mathcal{U}_{r+1}$ or $\mathcal{U}_{r}\subsetneq\mathcal{U}_{r+1}$.
The proof is by contradiction, that is, assume that (C.16) does not hold. Therefore, in view of the above observation, there exists $r\in\{1,\ldots,r^{*}-1\}$ such that $\mathcal{U}_{r}\subsetneq\mathcal{U}_{r+1}$. Then,
@@ 2. Vector $\mathbf{x}^{*}$ is a vertex allocation:
Hence (C.28) is proved.
If $\mathbf{x}^{*}$ is a regular allocation, in view of Remark C.1, the same reasoning leading to inequality (C.35) clearly remains valid for $r=r^{*}-1,\,r^{*}\geq 2$. ∎
@@
    where the fact that the above inequality is met for $h\in\mathcal{L}_{r^{*}-1}$ follows from (C.28). By comparing (C.45) and (C.46) we clearly see that (C.38) is satisfied. Lastly, equation (C.39) is fulfilled due to
@@
    where the first equality follows from (C.1) while the second one follows from Step 3 of the RNABOX after referring to (D.3) and using the fact that $\mathcal{U}_{r^{*}}=\mathcal{H}\setminus\mathcal{L}_{r^{*}}$ for a vertex allocation.
@@
$x_{h}^{*}=\begin{cases}M_{h},&\mathrlap{\qquad h\in\mathcal{U}^{*}}\\ A_{h}\,s(\emptyset,\,\mathcal{U}^{*}),&\mathrlap{\qquad h\in\mathcal{H}\setminus\mathcal{U}^{*},}\end{cases}$ (D.1)
@@ Appendix E Appendix: Convex optimization scheme and the KKT conditions
where $\mathbf{x}$ is the optimization variable, $\mathcal{D}\subseteq{\mathbb{R}}^{p},\,p\in{\mathbb{N}}_{+}$, the objective function $f:\mathcal{D}_{f}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}}$ and inequality constraint functions $g_{j}:\mathcal{D}_{g_{j}}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}},\,j=1,\ldots,\ell$, are convex, whilst equality constraint functions $w_{i}:\mathcal{D}_{w_{i}}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}},\,i=1,\ldots,k$, are affine. Here, $\mathcal{D}=\mathcal{D}_{f}\cap\bigcap_{i=1}^{k}\mathcal{D}_{w_{i}}\cap\bigcap_{j=1}^{\ell}\mathcal{D}_{g_{j}}$ denotes a common domain of all the functions. Point $\mathbf{x}\in\mathcal{D}$ is called feasible if it satisfies all of the constraints, otherwise the point is called infeasible. An optimization problem is called feasible if there exists $\mathbf{x}\in\mathcal{D}$ that is feasible, otherwise the problem is called infeasible.
@@
A point $\mathbf{x}^{*}\in\mathcal{D}\subseteq{\mathbb{R}}^{p},\,p\in{\mathbb{N}}_{+}$, is a solution to the convex optimization problem (E.1) in which functions $g_{j},\,j=1,\ldots,\ell$, are affine if and only if there exist numbers $\lambda_{i}\in{\mathbb{R}}$, $i=1,\ldots,k$, and $\mu_{j}\geq 0$, $j=1,\ldots,\ell$, called KKT multipliers, such that
$\begin{gathered}\nabla f(\mathbf{x}^{*})+\sum_{i=1}^{k}\lambda_{i}\nabla w_{i}(\mathbf{x}^{*})+\sum_{j=1}^{\ell}\mu_{j}\nabla g_{j}(\mathbf{x}^{*})=\mathbf{0}\\ w_{i}(\mathbf{x}^{*})=0,\mathrlap{\qquad i=1,\ldots,k}\\ g_{j}(\mathbf{x}^{*})\leq 0,\mathrlap{\qquad j=1,\ldots,\ell}\\ \mu_{j}g_{j}(\mathbf{x}^{*})=0,\mathrlap{\qquad j=1,\ldots,\ell.}\end{gathered}$ (E.2)